core: 0 of 1 pt 15 of 20 (19 complete) HW Score: 70%, 14 1.2.123 E Question He The chart shows the cost of tuition at a certain state university. Model the data in the chart with a linear function, using the points (1, 9989) and (3,11258) Predict the cost of college tuition in 2004-2005 College Year, x 997-1998,0 1998-1999, 1 1999-2000,2 2000-2001.3 2001-2002, 4 Estimated Tuition, y 59488 5 9989 $10583 $11258 $11972 What is the inear model for the data? (Type your answer in slope-intercept form. Use integers or decimals for any numbers in the expression Round to the nearest tenth as needed ) Enter your answer in the ansver bo r and then ck Chec Ansoe
Answer #1

Forming of a linear model when two coordinates are given.

First find slope and then find slope intercept form of a line.

In the end put the value of x to get the required result.

1, 9989 3, 11258) : 3, neu___ 1997-1998 1998-1979 2E0 yー9924 1258-9931--(x-1)- 3y-- 9989 = 634.5 C2-1) 634.5%+ q354-5 =그 634.5 (7) + 9 354,5 = 4441.5 ±9354.5 = $ 13796